3. sin 2x 3 4. sec(-210°) 5. 571 sin STE 6. CSC CE)

no cal
Answer #1

③ sin (-25) = -sin (26) 19 sin (St) = sin(x + -- sin(+3) = -sin (W) --CS(T) © csc (3A) - c(-210°) = Tz F-V3 2 sin (31) 4 sec I hope you have got your answer , if you need any further clarification in any step, please let me know through comment.